OpenStudy (anonymous):

Factor 6(2p + q)^2 -5(2p+q) - 25

OpenStudy (anonymous):

let 2p+q=x 6x^2-5x-25 6*-25=150 15-10=5 15*-10=150 make factors then substitute the value of x